OpenStudy (anonymous):

Solve the following \[\frac{3}{x+4} le 2\] express as interval and graph

OpenStudy (anonymous):

\[\frac{3}{x+4} \le 2\]

Parth (parthkohli):

First - the exception would be \(x \ne -4\) as getting -4 in the place of x would get you undefined in the fraction.

Parth (parthkohli):

Anyway, if you just kick out the exception — you can solve the inequality by multiplying x + 4 to both sides.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3 \le 2(x + 4)}\) \( \color{Black}{\Rightarrow 3 \le 2x + 8}\) \( \color{Black}{\Rightarrow 3 - 8 \le 2x}\) \( \color{Black}{\Rightarrow -5 \le 2x }\) \( \color{Black}{\Rightarrow -2.5 \le x}\)

OpenStudy (anonymous):

I thought you solve rational inequalities different...

Parth (parthkohli):

No, not really.

OpenStudy (anonymous):

\[3 \le 2x+8\] Solve for x..

OpenStudy (anonymous):

what if 2 was 0

Parth (parthkohli):

I don't know what your question is.

OpenStudy (anonymous):

Then no solution will be the answer..

Parth (parthkohli):

As I told you - the exception to x is -4.

OpenStudy (anonymous):

and 3 is not le or = 0...

Parth (parthkohli):

\( \color{Black}{\Rightarrow (-\infty , -5]\cup [-3,+\infty) }\) This is the interval notation.

myininaya (myininaya):

\[\frac{3}{x+4}-2 \le 0\] Figure out when the expression on the left is 0 and when it is undefeined \[\frac{3-2(x+4)}{x+4} \le 0\] \[\frac{3-2x-8}{x+4} \le 0\] \[\frac{-2x-5}{x+4} \le 0\]

myininaya (myininaya):

Then after you figure out when that expression of the left is undefined and zero you will test the intervals around each number to see when you have less than 0

OpenStudy (anonymous):

so were only concerned about the denominator @myininaya

myininaya (myininaya):

no that is not what i said

myininaya (myininaya):

when the top is 0 and when the expression is undefined

myininaya (myininaya):

when -2x-5=0 x+4=0

Parth (parthkohli):

No! The bottom - not the top!

myininaya (myininaya):

test around those numbers

OpenStudy (anonymous):

:\ :{ :[

myininaya (myininaya):

you already know -2x-5=0 will give you one value of x that satisfies the inequality because your expression also includes the equals part

OpenStudy (anonymous):

Is the interval [-4,-5/2]

OpenStudy (anonymous):

x cannot be -4...

Parth (parthkohli):

I told you the interval notation @purplec16

OpenStudy (anonymous):

I don't see how you got that @ParthKohli

OpenStudy (anonymous):

(-4,5/2]

myininaya (myininaya):

-2x-5=0 when x=5/-2=-5/2 x+4=0 when x=-4 We already know -5/2 is part of the solution while -4 is not but we use these numbers to test around -----|----|---- -4 -5/2

OpenStudy (anonymous):

test outside the interval?

myininaya (myininaya):

Plug in some number before -4 Plug in some number btw -4 and -5/2 Plug in some number after -5/2 And if the number you get is less than 0 then that interval is part of the solution

OpenStudy (anonymous):

test the whole domain...

OpenStudy (anonymous):

I did that... so is my answer correct?

myininaya (myininaya):

\[\frac{-2x-5}{x+4} \le 0\] A number before -4: I will do -100 for fun! :) So \[\frac{-2(-100)-5}{-100+4}=\frac{200-5}{-96}=\frac{195}{-96} <0\] This interval works Now you try the second interval and third interval (we only have three intervals to look at )

OpenStudy (anonymous):

I still don't see the answer? (-infity, -4)U(-4,5/2]U[5/2,infity)

myininaya (myininaya):

No one ever gave you the answer We are trying to help you get the answer on your own Do you not understand what I'm asking you?

OpenStudy (anonymous):

nope...

myininaya (myininaya):

to do*

myininaya (myininaya):

Ok I tried the first interval do you see that?

OpenStudy (anonymous):

I did what you ask and that's what I got, it's not correct?

myininaya (myininaya):

why do you have 5/2 in there?

myininaya (myininaya):

-5/2 was the zero

myininaya (myininaya):

This is what I said to do Plug in some number before -4 Plug in some number btw -4 and -5/2 Plug in some number after -5/2 And if the number you get is less than 0 then that interval is part of the solution

OpenStudy (anonymous):

(-infity, -4)U(-4,-5/2]U[-5/2,infity)

myininaya (myininaya):

Then I did the first interval which were the numbers before -4 for you

myininaya (myininaya):

darn it! I got the aw snap page :( Okay yeah that works!

OpenStudy (anonymous):

So am right :D

OpenStudy (anonymous):

So am right :D?

myininaya (myininaya):

no i'm sorry try -3 what did you get i was talking about the page sorry

myininaya (myininaya):

One of your intervals is wrong is what I'm saying

myininaya (myininaya):

if you plug in a value btw -4 and -5/2 what do you get ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

0

myininaya (myininaya):

no 1 is right not 0 that is bigger than zero so that does not work with our inequality

myininaya (myininaya):

so what can we say about that interval?

OpenStudy (anonymous):

well undefined...

OpenStudy (anonymous):

0 for -5/2 and undefined for -4

myininaya (myininaya):

no 1 you plug in -3 you get 1 not 0 and not undefined

myininaya (myininaya):

no....

myininaya (myininaya):

i said plug in a number btw those numbers

OpenStudy (anonymous):

I really don't know what to say about this interval...

OpenStudy (anonymous):

\[(-\infty,-4)U(-\frac{5}{2},\infty)\]

myininaya (myininaya):

ok please try to do what i say: Plug in a number between -4 and -5/2 into the expression \[\frac{-2x-5}{x+4}\] We are looking at the middle interval right now We are looking at (-4,-5/2] to see if it should be included in the answer I picked a number for you that number being -3 Now if you plug in that number -3 into \frac{-2x-5}{x+4}\] You will get what you said which was 1 and 1 is bigger than 0 not less than 0 So we do not include the second interval make sure you include your zeros of \frac{-2x-5}{x+4}\] Remember x=-5/2 was a zero so it should be included Your answer is lacking a little